# Dispersion of first sound in a weakly interacting ultracold Fermi liquid: an exact calculation

### v2

@@ Acknowledgements.
H.K. acknowledges support from the EUR grant NanoX n°ANR-17-EURE-0009 in the framework of the “Programme des Investissements d’Avenir”. N.N acknowledges support by the AFOSR (Grant No. FA9550-23-1-0605), NSF (Grant No. PHY-1945324), DARPA (Grant No. HR00112320038), and the David and Lucile Packard Foundation. T.R. and H.K. thank Yale University for its hospitality.
.1 Observability of the spectrum in the density response function
@@
On the hydrodynamic side of the crossover ($\omega_{0}\tau<2$), and for $k_{\textrm{F}}a=0$ ($c_{1}=v_{\textrm{F}}/\sqrt{3}$) we fit $\text{Im}\bar{\chi}_{\rho}(c+\textrm{i}0^{+})$ to a Lorentzian function:
$f_{\rho}^{\rm fit}(c)=\frac{Z}{(c-\bar{c}_{1}-B\omega_{0}^{2}\tau^{2})^{2}+A^{2}\omega_{0}^{2}\tau^{2}}$ (33)
@@
$A$ $\underset{\omega_{0}\tau\to 0}{\to}$ $\frac{\textrm{Re}(z_{1}-\bar{c}_{1})}{(\omega_{0}\tau)^{2}}=\frac{2\sqrt{3}\pi^{2}}{225}\frac{5t_{\eta}^{2}-\tau_{\eta}^{2}}{\tau^{2}}$ (34)
$B$ $\underset{\omega_{0}\tau\to 0}{\to}$ $-\frac{\textrm{Im}(z_{1})}{\omega_{0}\tau}=\frac{2\pi}{15}\frac{\tau_{\eta}}{\tau}$ (35)
Figure 2: The transition from hydrodynamic (red curves) to collisionless regime (blue curves) in the density response of a weakly interacting Fermi gas with $k_{\textrm{F}}a=-0.5$. The broadening and the positive shift of the resonance frequency from the first sound velocity (vertical dashed line, Eq. (24)) is clearly visible at small $\omega_{0}\tau$. The curve are obtained by numerically solving Eq. (19) truncated to $n_{\rm max}=100$ and $l_{\rm max}$ ranging from 20 in the hydrodynamic regime to 8500 in the collisionless regime.
Figure 3: The sound attenuation and dispersion coefficients $\gamma$ and $\Delta c$ obtained by fitting a Lorentzian function Eq. (33) (green and orange lines in inset, respectively for $\omega_{0}\tau=0.04$ and $\omega_{0}\tau=2$) to the density response $\bar{\chi}_{\rho}$ obtained by numerically solving Eq. (19) at non vanishing $\omega_{0}\tau$. The horizontal dashed and dotted line show respectively the exact hydrodynamic limits Eqs. (34) and (35), and the relaxation time approximation ($\tau_{\eta},t_{\eta}\to\tau_{\eta}^{(0)}$). This figure is drawn in the limit $k_{\textrm{F}}a\to 0$ ($c_{1}/v_{\textrm{F}}=1/\sqrt{3}$).

### v3

@@ Acknowledgements.
H.K. acknowledges support from the French Agence Nationale de la Recherche (ANR), under grant ANR-23-ERCS-0005 (project DYFERCO) and the EUR grant NanoX n°ANR-17-EURE-0009 in the framework of the “Programme des Investissements d’Avenir”. N.N acknowledges support by the AFOSR (Grant No. FA9550-23-1-0605), NSF (Grant No. PHY-1945324), DARPA (Grant No. HR00112320038), and the David and Lucile Packard Foundation. T.R. and H.K. thank Yale University for its hospitality.
Appendix A Collsion kernel
We write the linearized collision integral as
$I_{\rm lin}=I_{\rm lin,\uparrow}+I_{\rm lin,\downarrow}=-\Gamma(\textbf{p})\delta n(\textbf{p})\\ -\frac{1}{L^{3}}\sum_{\textbf{p}^{\prime}}\left[E(\textbf{p}^{\prime},\textbf{p})-2S(\textbf{p}^{\prime},\textbf{p})\right]\delta n(\textbf{p}^{\prime})$ (44)
with the collision kernels $E$ and $S$ (notice that we swap the arguments $\textbf{p}\leftrightarrow\textbf{p}^{\prime}$)
$E(\textbf{p},\textbf{p}^{\prime})$ $=$ $\frac{2\pi g^{2}}{L^{3}}\sum_{\textbf{p}_{3},\textbf{p}_{4}}\delta^{\textbf{p}+\textbf{p}^{\prime}}_{\textbf{p}_{3}+\textbf{p}_{4}}\delta\left(\epsilon_{\textbf{p}}+\epsilon_{\textbf{p}^{\prime}}-\epsilon_{\textbf{p}_{3}}-\epsilon_{\textbf{p}_{4}}\right)\left[n_{\rm eq}(\epsilon_{\textbf{p}^{\prime}})\overline{n}_{\rm eq}(\epsilon_{\textbf{p}_{3}})\overline{n}_{\rm eq}(\epsilon_{\textbf{p}_{4}})+\overline{n}_{\rm eq}(\epsilon_{\textbf{p}^{\prime}})n_{\rm eq}(\epsilon_{\textbf{p}_{3}})n_{\rm eq}(\epsilon_{\textbf{p}_{4}})\right]$ (45)
$S(\textbf{p},\textbf{p}^{\prime})$ $=$ $\frac{2\pi g^{2}}{L^{3}}\sum_{\textbf{p}_{2},\textbf{p}_{4}}\delta^{\textbf{p}+\textbf{p}_{2}}_{\textbf{p}^{\prime}+\textbf{p}_{4}}\delta\left(\epsilon_{\textbf{p}}+\epsilon_{\textbf{p}_{2}}-\epsilon_{\textbf{p}^{\prime}}-\epsilon_{\textbf{p}_{4}}\right)\left[n_{\rm eq}(\epsilon_{\textbf{p}_{2}})\overline{n}_{\rm eq}(\epsilon_{\textbf{p}^{\prime}})\overline{n}_{\rm eq}(\epsilon_{\textbf{p}_{4}})+\overline{n}_{\rm eq}(\epsilon_{\textbf{p}_{2}})n_{\rm eq}(\epsilon_{\textbf{p}^{\prime}})n_{\rm eq}(\epsilon_{\textbf{p}_{4}})\right]$ (46)
with $\bar{n}_{\rm eq}=1-n_{\rm eq}$, and the short-hand notation $\epsilon_{\textbf{p}^{\prime\prime}}$ for $\epsilon_{\rm 0}(\textbf{p}^{\prime\prime})$. The quasiparticule lifetime $\Gamma$ is related to $E$ and $S$ by $\Gamma(\textbf{p})=\sum_{\textbf{p}^{\prime}}E(\textbf{p},\textbf{p}^{\prime})/L^{3}=\sum_{\textbf{p}^{\prime}}S(\textbf{p},\textbf{p}^{\prime})/L^{3}$, and the total collision kernel obeys the conservation laws
$\Gamma(\textbf{p})\textbf{p}^{k}+\frac{1}{L^{3}}\sum_{\textbf{p}^{\prime}}\left[E(\textbf{p}^{\prime},\textbf{p})-2S(\textbf{p}^{\prime},\textbf{p})\right]\textbf{p}^{\prime k}=0,\ k=0,1,2$ (47)
Using the property of thermal equilibrium
$n_{\rm eq}(\epsilon_{\textbf{p}_{1}})n_{\rm eq}(\epsilon_{\textbf{p}_{2}})\overline{n}_{\rm eq}(\epsilon_{\textbf{p}_{3}})\overline{n}_{\rm eq}(\epsilon_{\textbf{p}_{4}})\\ =\overline{n}_{\rm eq}(\epsilon_{\textbf{p}_{1}})\overline{n}_{\rm eq}(\epsilon_{\textbf{p}_{2}}){n}_{\rm eq}(\epsilon_{\textbf{p}_{3}}){n}_{\rm eq}(\epsilon_{\textbf{p}_{4}})$ (48)
(for four momenta on the energy shell: $\epsilon_{\textbf{p}_{1}}+\epsilon_{\textbf{p}_{2}}=\epsilon_{\textbf{p}_{3}}+\epsilon_{\textbf{p}_{4}}$) together with the change of variable Eq. (6) and the properties
$E(\textbf{p}^{\prime},\textbf{p})\frac{n_{\rm eq}(\epsilon_{\textbf{p}^{\prime}})\bar{n}_{\rm eq}(\epsilon_{\textbf{p}^{\prime}})}{n_{\rm eq}(\epsilon_{\textbf{p}})\bar{n}_{\rm eq}(\epsilon_{\textbf{p}})}$ $=$ $E(\textbf{p},\textbf{p}^{\prime})$ (49)
$S(\textbf{p}^{\prime},\textbf{p})\frac{n_{\rm eq}(\epsilon_{\textbf{p}^{\prime}})\bar{n}_{\rm eq}(\epsilon_{\textbf{p}^{\prime}})}{n_{\rm eq}(\epsilon_{\textbf{p}})\bar{n}_{\rm eq}(\epsilon_{\textbf{p}})}$ $=$ $S(\textbf{p},\textbf{p}^{\prime})$ (50)
$T\frac{\partial n_{\rm eq}}{\partial\epsilon_{0}}\Big{|}_{\epsilon_{0}=\epsilon_{\textbf{p}}}=-n_{\rm eq}(\epsilon_{\textbf{p}})(1-n_{\rm eq}(\epsilon_{\textbf{p}}))$ (51)
we swap the arguments of $E$ and $S$ as we move the transport equation Eq. (5) from the distribution $\delta n$ to $\nu$:
$\left(\omega-\frac{\textbf{p}\cdot\textbf{q}}{m}\right)\nu(\textbf{p})+\frac{\textbf{p}\cdot\textbf{q}}{m}\left(g\delta\rho+U\right)\\ =-\textrm{i}\left(\Gamma(\textbf{p})\nu(\textbf{p})+\frac{1}{L^{3}}\sum_{\textbf{p}^{\prime}}\left[E(\textbf{p},\textbf{p}^{\prime})-2S(\textbf{p},\textbf{p}^{\prime})\right]\nu(\textbf{p}^{\prime})\right)$ (52)
Note that in terms of $\nu$, the fluctuations of the total density are $\delta\rho=(1/V)\sum_{\textbf{p}^{\prime}}({\partial n_{\rm eq}}/{\partial\epsilon_{\rm 0}})_{\epsilon_{\rm 0}=\epsilon_{\textbf{p}^{\prime}}}\nu(\textbf{p}^{\prime})$.
In Eqs. (45)–(46), we eliminate $\textbf{p}_{4}$ using momentum conservation, and convert the sums into integrals using
$\frac{1}{L^{3}}\sum_{\textbf{p}_{i}}\to\frac{mTp_{\rm F}}{(2\pi)^{2}}\int_{-\infty}^{+\infty}\textrm{d}\epsilon_{i}\int_{0}^{\pi}\sin\vartheta_{i}\textrm{d}\vartheta_{i}\int_{0}^{2\pi}\frac{\textrm{d}\varphi_{i}}{2\pi}$ (53)
where $i=2,3$ for $E$ and $S$ respectively, $\epsilon_{i}=(\epsilon_{\rm 0}(\textbf{p}_{i})-\mu)/T$, $\theta_{i}$ is the polar angle between $\textbf{p}_{3}$ and $\textbf{p}+\textbf{p}^{\prime}$ or $\textbf{p}_{2}$ and $\textbf{p}-\textbf{p}^{\prime}$. Such choices of spherical frames, together with the isotropic collision probability $g^{2}$, guarantees that the integrand is independant of the azimuthal angles $\varphi_{i}$. Note that the conversion Eq. (53) approximates $p_{i}^{2}\textrm{d}p_{i}\simeq mTp_{F}\textrm{d}\epsilon_{i}$, which is valid up to corrections in $O(T)$. To leading order in $T$, the energy-conservation constraints can be replaced by their $T=0$ version, namely
$\!\!\!\!\!\!\!\!\!\epsilon_{\textbf{p}}+\epsilon_{\textbf{p}^{\prime}}-\epsilon_{\textbf{p}_{3}}-\epsilon_{\textbf{p}_{4}}\!\!\!$ $=$ $\!\!\!\frac{2p_{\rm F}^{2}}{m}\left[\cos\frac{\alpha}{2}\left(\cos\theta_{3}\!-\!\cos\frac{\alpha}{2}\right)\right]$ (54)
$\!\!\!\!\!\!\!\!\!\epsilon_{\textbf{p}}+\epsilon_{\textbf{p}_{2}}-\epsilon_{\textbf{p}^{\prime}}-\epsilon_{\textbf{p}_{4}}\!\!\!$ $=$ $\!\!\!\frac{2p_{\rm F}^{2}}{m}\left[\sin\frac{\alpha}{2}\left(\cos\theta_{2}\!-\!\sin\frac{\alpha}{2}\right)\right]$ (55)
The Dirac function can then be used to perform the remaining angular integration over $\theta_{i}$. There remains to integrate over the reduced energy $\epsilon_{i}$, using:
$\int_{-\infty}^{\infty}\textrm{d}\epsilon_{3}\left[n(\epsilon^{\prime})\left(1-n(\epsilon_{3})-n(\epsilon+\epsilon^{\prime}-\epsilon_{3})\right)+n(\epsilon_{3})n(\epsilon+\epsilon^{\prime}-\epsilon_{3})\right]$ $=$ $\mathcal{S}(\epsilon,-\epsilon^{\prime})$ (56)
$\int_{-\infty}^{\infty}\textrm{d}\epsilon_{2}\left[n(\epsilon_{2})\left(1-n(\epsilon^{\prime})-n(\epsilon+\epsilon_{2}-\epsilon^{\prime})\right)+n(\epsilon^{\prime})n(\epsilon+\epsilon_{2}-\epsilon^{\prime})\right]$ $=$ $\mathcal{S}(\epsilon,\epsilon^{\prime})$ (57)
The result are collision kernels whose angular and energy dependance are factorized:
$E(\textbf{p},\textbf{p}^{\prime})$ $=$ $\frac{m^{2}T}{2\pi p_{\rm F}}\frac{g^{2}}{2\cos(\theta/2)}\mathcal{S}(\epsilon,-\epsilon^{\prime})$ (58)
$S(\textbf{p},\textbf{p}^{\prime})$ $=$ $\frac{m^{2}T}{2\pi p_{\rm F}}\frac{g^{2}}{2\sin(\theta/2)}\mathcal{S}(\epsilon,\epsilon^{\prime})$ (59)
This remarkable property is specific to Fermi liquids and distinguishes them from classical gases. It is a direct consequence of the freezing of the collisions to the vicinity of the Fermi surface when $T\to 0$.
To derive the transport equation in the low-temperature limit Eq. (7), there remains to divide Eq. (5) by $\omega_{0}(\partial n_{\rm eq}/\partial\epsilon_{0})_{\epsilon_{0}=\epsilon_{0}(\textbf{p})}$ and expand each term in powers of $T$ at fixed $\epsilon$ and $\epsilon^{\prime}$ (which amounts to replacing $p$ and $p^{\prime}$ by $p_{\rm F}$ outside functions that are peaked about the Fermi sphere).
Appendix B Observability of the spectrum in the density response function
@@
On the hydrodynamic side of the crossover ($\omega_{0}\tau_{\sigma}<0.5$), and for $k_{\textrm{F}}a=0$ ($c_{1}=v_{\textrm{F}}/\sqrt{3}$) we fit $\text{Im}\bar{\chi}_{\rho}(c+\textrm{i}0^{+})$ to a Lorentzian function:
$f_{\rho}^{\rm fit}(c)=\frac{Z}{(c-\bar{c}_{1}-B\omega_{0}^{2}\tau_{\sigma}^{2})^{2}+A^{2}\omega_{0}^{2}\tau_{\sigma}^{2}}$ (60)
@@
$B$ $\underset{\omega_{0}\tau\to 0}{\to}$ $\frac{\textrm{Re}(z_{1}-\bar{c}_{1})}{(\omega_{0}\tau_{\sigma})^{2}}=\frac{2\sqrt{3}}{225}\frac{5t_{\eta}^{2}-\tau_{\eta}^{2}}{\tau_{\sigma}^{2}}$ (61)
$A$ $\underset{\omega_{0}\tau\to 0}{\to}$ $-\frac{\textrm{Im}(z_{1})}{\omega_{0}\tau_{\sigma}}=\frac{2}{15}\frac{\tau_{\eta}}{\tau_{\sigma}}$ (62)
To estimate the order of magnitude of the second-order hydrodynamical effects in experiments, we use the parameters $k_{\rm F}a=-0.67$, $T/T_{\rm F}=0.17$, $q/k_{\rm F}=0.01$ taken from Ref. [21] (see Fig. 4 therein). This corresponds to $\omega_{0}\tau_{\sigma}\simeq 0.42$ for which a first sound resonance is still clearly visible in Fig. 4a of Ref. [21]. Plugging these experimental parameters in Eq. (37), we estimate the relative frequency shift and damping rate to be
$\frac{\omega_{q}}{c_{1}q}-1$ $\simeq$ $3\%$ (63)
$\frac{\Gamma_{q}}{c_{1}q}$ $\simeq$ $13\%$ (64)
The small value of $\omega_{q}-c_{1}q$ is challenging to measure, but it could be within experimental reach in the near future. This would allow to distinguish between supersonic $\omega_{q}>c_{1}q$ and subsonic $\omega_{q}<c_{1}q$ dispersion.
Figure 2: The transition from hydrodynamic (red curves) to collisionless regime (blue curves) in the density response of a weakly interacting Fermi gas with $k_{\textrm{F}}a=-0.5$. The broadening and the positive shift of the resonance frequency from the first sound velocity (vertical dashed line, Eq. (32)) is clearly visible at small $\omega_{0}\tau_{\sigma}$. The curve are obtained by numerically solving Eq. (21) truncated to $n_{\rm max}=100$ and $l_{\rm max}$ ranging from 20 in the hydrodynamic regime to 8500 in the collisionless regime.
Figure 3: The sound attenuation and dispersion coefficients $\gamma$ and $\Delta c$ obtained by fitting a Lorentzian function Eq. (60) (green and orange lines in inset, respectively for $\omega_{0}\tau_{\sigma}=0.01$ and $\omega_{0}\tau_{\sigma}=0.5$) to the density response $\bar{\chi}_{\rho}$ obtained by numerically solving Eq. (21) at non vanishing $\omega_{0}\tau_{\sigma}$ (symbols in the inset). The horizontal dashed and dotted line show respectively the exact hydrodynamic limits Eqs. (61) and (62), and the relaxation time approximation ($\tau_{\eta},t_{\eta}\to\tau_{\eta}^{(0)}$). This figure is drawn in the limit $k_{\textrm{F}}a\to 0$ ($c_{1}/v_{\textrm{F}}=1/\sqrt{3}$).